An ion M+a has the magnetic moment equal to √24 BM. the value of a is(AT.no of M =24)

Grade:11

1 Answers

jjijijijij
5 years ago
we know that the formula for magnatic moment=\sqrt{n(n+2)} where n is no. of unbonded electrons.
as the ion has a charge of +a then a=no.of unbonded electrons.
\sqrt{a(a+2)}=\sqrt{24}BM
squaring both sides
a(a+2)=24
a2+2a-24=0
a2+6a-4a-24=0
a(a+6)-4(a+6)
we conclude that a has 2 values -6 and + 4
since M has a positive charge we may have the result as a =4